# Searching for Exotrojans in Pulsar Binary Systems

### v1

@@ Abstract
Trojan asteroids are found in the equilateral triangle Lagrange points of the Sun-Jupiter system in great number, though they also exist less prolifically in other Sun-planet systems. Despite up to planetary mass Trojans being predicted in extrasolar systems (i.e. exotrojans), they remain largely unconfirmed, though with recent strong candidate evidence emerging. We turn the current search for exotrojans to radio pulsars with low-mass companions ($\sim 0.01\,\rm{M}_{\odot}$) using accurately measured pulse times of arrival. With techniques developed for detecting the reflex motion of a star due to a librating Trojan, we place reasonable mass constraints ($\sim 1\,\rm{M}_{\oplus}$) on potential exotrojans around binary pulsars observed in the NANOGrav 15-year data set. We find weak evidence consistent with $\sim 1\,\rm{M}_{\rm J}$ exotrojans in the PSR J0023+0923 and PSR J1705-1903 systems, though the signals likely have a different, unknown source. We also place a libration-independent upper mass constraint of $\sim 8$ MJ on exotrojans in the PSR 1641+8049 binary system by looking for an inconsistency between the times of superior conjunction as measured by optical light curves and those predicted by radio timing.
@@
facilities: GTC, GBT, Arecibosoftware: PINT (J. Luo et al., 2019, 2021), rebound (H. Rein & S. F. Liu, 2012; H. Rein & D. S. Spiegel, 2015), ENTERPRISE (J. A. Ellis et al., 2019; A. D. Johnson et al., 2023), enterprise_extensions (S. R. Taylor et al., 2021; A. D. Johnson et al., 2023), PTMCMCSampler (J. Ellis & R. van Haasteren, 2017; A. D. Johnson et al., 2023), numpy (C. R. Harris et al., 2020), matplotlib (J. D. Hunter, 2007), corner (D. Foreman-Mackey, 2016).
1 Introduction
Trojans are objects that reside in the $L_{4}$ or $L_{5}$ Lagrange points of two massive bodies. These so-called co-orbitals share the same mean motion as the more-massive companion (known as the dominant companion) around its higher-mass host, and they appear to trail or lead the companion by $60^{\circ}$ throughout its orbit. In the Jupiter-Sun system alone, there are over ten thousand confirmed Trojans (e.g., J. Li et al., 2023) along with estimates of hundreds of thousands more that have diameters above $\sim$ 2 km (see, e.g., D. C. Jewitt et al., 2000; F. Yoshida & T. Nakamura, 2005). Every Solar System planet has Trojans of its own, barring Mercury—which may lack confirmed Trojans due to observational difficulties—and there are even Trojans of the Saturn-Tethys system.
The abundance of Trojans in our Solar System suggests that extrasolar Trojans (exotrojans) should also be common. Indeed, any robust Solar System formation theory must allow for their creation, and many of these theories do not just predict Trojan asteroids but also, surprisingly, Trojans of planetary mass (e.g., G. Laughlin & J. E. Chambers, 2002; P. Cresswell & R. P. Nelson, 2008; W. Lyra et al., 2009; M. Montesinos et al., 2020). Once formed, it has been shown that these Trojans can find themselves in relatively large stable regions, as detailed in J. M. A. Danby (1964); C. Giuppone et al. (2010); A. Leleu et al. (2015). In the case of circular orbits, the triangular Lagrange points are only stable if
@@ 1 Introduction
where $m_{\rm p}$ is the mass of the host star (a pulsar in this study), $m_{\rm c}$ is the mass of the star’s dominant companion, and $m_{\rm T}$ is the mass of the Trojan (M. Gascheau, 1843). J. M. A. Danby (1964) showed that introducing a non-zero orbital eccentricity generally decreases the window of stable masses.
@@
Unless the Trojan is librating around its stable point, the gravitational effects from a single orbiting companion are degenerate with those of two co-orbiting companions up to $\mathcal{O}(e_{\rm{o}}^{2})$ (A. Leleu et al., 2017), where $e_{\rm o}$ is the orbital eccentricity. This degeneracy helps explain why, despite their prevalence in our Solar System and their accompanying formation models, no Trojans outside the Solar System have yet been confirmed. E. B. Ford & B. S. Gaudi (2006) showed that combining radial velocity data with transit-timing data breaks this degeneracy. A. Leleu et al. (2017) further formalized their approach into the robust $\alpha$-test method to include the results of several transits. O. Balsalobre-Ruza et al. (2024) applied this technique to identify one of the strongest candidates to date with a $3$-$\sigma$ detection, building on the earlier works of J. Lillo-Box et al. (2018a, b). These works are part of, or helped form the foundation for, the broader TROY project,111www.troy-project.com which seeks to discover the first extrasolar Trojan planets. While most of their work has been confined to optical observations of main-sequence stars, O. Balsalobre-Ruza et al. (2023) potentially identified a forming Trojan gas & dust accumulation in the PDS 70 protoplanetary disk (e.g., S. Haffert et al., 2019).
@@
Other studies have searched for photometric transits of the Trojans themselves. M. Janson (2013) used the Kepler space telescope to rule out transiting Trojans down to 1 R in 2244 systems, and R. Moldovan et al. (2010) used the MOST satellite to place a lunar mass upper limit on a Trojan in the HD 209458 system. Rather than looking for transits from individual trojans, M. Hippke & D. Angerhausen (2015) looked for a cumulative Trojan effect across nearly 4000 planets by adding their Kepler light curves coherently by orbital phase, and identified potential transits of the stacked light curves at the expected $L_{4}$ and $L_{5}$ orbital phases. The main problem with identifying Trojans transiting their host stars is the fact that we expect exotrojans to have a broad range of orbital inclinations, as is the case for the Jupiter Trojans (F. Yoshida & T. Nakamura, 2005; D. Nesvornỳ et al., 2013; W. F. Bottke et al., 2023). Therefore, the dominant Kepler planet transiting its host star does not guarantee that any Trojan planets will also transit during any given observation. In contrast, the $\alpha$-test method does not rely on in-plane Trojans.
@@
While almost every exotrojan search has been done around main-sequence stars, as detailed above, no study has looked for them around pulsar binary systems. Just as we expect non-pulsar Trojans to form around ordinary stars during the protoplanetary disk phase (e.g., M. Montesinos et al., 2020), Trojans may also form around pulsars when a disk is present and then settle via gas drag into stable so-called tadpole (T-type) orbits (W. Lyra et al., 2009). Such a disk can form in at least two possible ways. One type of pulsar disk, the supernova fallback disk, forms when some of the material from the supernova that formed the pulsar returns to the pulsar and collapses into a disk (E. A. Behrens et al., 2020; K. Xu et al., 2024). Such disks are a plausible formation channel for pulsar planets (D. Lin et al., 1991; A. Wolszczan & D. A. Frail, 1992), and would explain why the planets around PSR B1257+12 have similar inclinations (M. Konacki & A. Wolszczan, 2003). The other possible type of pulsar disk forms when the compact pulsar strips material from its evolving, ablated companion. This material forms an accretion disk, which collects onto the pulsar and transitions the pulsar into an MSP (A. M. Archibald et al., 2009; A. Papitto & D. d. Martino, 2021). Before the accretion disk dissipates, it may form planets (E. A. Behrens et al., 2020), an asteroid belt (R. Shannon et al., 2013), or Trojan planets in 1:1 resonance with the pulsar’s companion, as in the case of protoplanetary-disk-formed Trojans (W. Lyra et al., 2009; M. Montesinos et al., 2020). Trojans, too, may be stripped of their material, but the timescale of complete dissipation is unclear, with some estimates ranging up to a few Gyr for the dominant companion (e.g. S. Ginzburg & E. Quataert, 2020). This picture is complicated by the fact that material from the dominant companion could continue to fall into the Triangular Lagrange points. In any case, detecting an interloping Trojan may be possible even if it lasts just a few hundred Myr before complete dissipation.
@@
Ever since the discovery of the first exoplanet (A. Wolszczan & D. A. Frail, 1992), there has been a growing interest in discovering massive bodies in the environs of pulsars. Besides studies looking for planets in the ordinary P-type orbit (e.g., M. Bailes et al., 2011; R. Spiewak et al., 2018; E. A. Behrens et al., 2020; I. C. Niţu et al., 2022), pulsar asteroid belts (R. Shannon et al., 2013; R. J. Jennings et al., 2020a) and freely-floating planets gravitationally scattering with pulsars (R. J. Jennings et al., 2020b; L. Dey et al., 2025) have also been investigated. In this work, we turn the search to Trojans that may reside in black widow pulsar systems—binaries characterized by a millisecond pulsar (MSP) with a low-mass ($\sim 0.01\,\mathrm{M_{\odot}}$) companion (e.g., T. Shahbaz et al., 2017)—and are therefore likely to follow the stability criterion given in Equation 1.
@@
In order to conduct the first-ever search for Trojan planets in pulsar binary systems, we use the North American Nanohertz Observatory for Gravitational Waves (NANOGrav) 15-year (NG15yr) narrow-band dataset (G. Agazie et al., 2023) as well as optical-band light curves of PSR J1641+8049 from A. Y. Kirichenko et al. (2024) to search for pulsar systems bearing trojans in the T-type configuration. We do not search for the same in the horseshoe configuration because these systems do not yield a closed-form waveform in the pulse times of arrival (TOAs; see Section 3.2 and A. Leleu et al. 2015). A Trojan in the T-type configuration librates around a particular triangular Lagrange point, while a Trojan in the horseshoe configuration will repeatedly librate through $L_{4}$-$L_{3}$-$L_{5}$ Lagrange points (in that order) and back again. Excluding horseshoe orbiters does not greatly limit our search, as not only are they less common in the Solar System (e.g., M. Kaplan & S. Cengiz, 2020), but also their stability criterion are stricter (e.g., G. Laughlin & J. E. Chambers, 2002).
All uncertainties reported in this work enclose the 68% credible interval. We quote our upper limits at 95% credibility. We use both the parentheses-on-the-last-digit(s) convention and the $\pm$ notation to express these 68% uncertainties. Throughout, $\log$ denotes the base-10 logarithm.
2 Selected Pulsars and their Observational Data
According to the ATNF Catalog (R. N. Manchester et al., 2005), [^2] https://www.atnf.csiro.au/research/pulsar/psrcat there are 72 pulsar binary systems that satisfy Equation 1 when $m_{\rm T}=0$, $m_{\rm p}=1.35$, and $m_{\rm c}$ is determined by the Newtonian mass function (Equation 17) when $\sin{I}=1$, where $I$ is the pulsar system’s orbital inclination with respect to the plane of the sky. Out of these 72 systems, we restrict this study to PSR J1641+8049 and the eight systems that are in the NG15yr dataset (see Table 1) due to their high observation cadence, long timing baseline, and availability of the dataset. Other datasets also satisfy these criteria (e.g., J. Antoniadis et al., 2023; A. Zic et al., 2023), but we leave those analyses for future work due to differences in the noise modeling and observatories. For PSR J1641+8049, we use the method developed by E. B. Ford & B. S. Gaudi (2006) to seek out Trojan-caused inconsistencies between the pulsar-companion conjunction times as measured by optical lightcurves of the system (A. Y. Kirichenko et al., 2024) and those predicted by radio pulsar timing (see Section 3.1). For the eight NG15yr pulsars of interest, we adapt the method of A. Leleu et al. (2015) to look for TOA residuals caused by Trojan librations—a method that only requires TOAs (see Section 3.2).
2.1 Optical Data
The first optical observations of the PSR J1641+8049 counterpart were reported by R. S. Lynch et al. (2018), revealing strong brightness variations tied to the binary period. The detection triggered further optical studies: using the High PERformance CAMera (HiPERCAM) installed at the 10.4m Gran Telescopio Canarias (GTC), D. Mata Sánchez et al. (2023) performed phase-resolved multi-band photometric observations and analysed the light curves of the system, yielding its binary parameters and the distance. A. Y. Kirichenko et al. (2024) carried out independent phase-resolved multi-band optical studies with the Optical System for Imaging and low Resolution Integrated Spectroscopy (OSIRIS) at the GTC and performed a comparative light-curve analysis using both the HiPERCAM and OSIRIS datasets.
As demonstrated by both studies, the J1641+8049 light curves show a single-peaked sinusoidal-like modulation per orbital period with a maximum when the side of the companion heated by the neutron star is directed toward the observer (superior conjunction) and a minimum when the cold back side is visible (inferior conjunction). The system parameters obtained by A. Y. Kirichenko et al. (2024) through the light curve analysis, including the distance to the source 4.6–4.8 kpc and the orbital inclination 56–59, are found to be consistent with those of D. Mata Sánchez et al. (2023). However, the optical flux of the source at the maximum brightness phase faded by a factor of $\sim$2 as compared with the values reported in the D. Mata Sánchez et al. (2023) study. Despite this variability, the companion’s irradiated hemisphere remains among the hottest (8000–9500 K) observed in the known black widow systems.
2.2 Radio Pulsar Timing Data
The times of arrival for PSR 1641+8049 came from the CHIME/Pulsar backend (M. Amiri et al., 2021) on the Canadian Hydrogen Intensity Mapping Experiment (CHIME) radio telescope (M. Amiri et al., 2022). These were used in calculating the times of superior conjunction.
Of the eight selected NG15yr pulsars, four were observed with the Arecibo Observatory (PSRs J0023+0923, J1745+1017, J2212+3000, and J2234+0944) and four with the Green Bank Telescope (PSRs J0610-2100, J0636+5128, J1705-1903, and J1719-1438). The TOAs of our selected pulsars cover time spans ranging from 3.4 yr (PSR J0610-2100) to 9.0 yr (PSR J0023+0923). The near-random telescope cadences with respect to libration phase (see Equations 3.2 and 6), combined with the years-long timing baselines, ensure that the TOAs effectively sample all libration phases. Therefore, the Nyquist–Shannon sampling theorem criterion is relaxed, and the libration frequencies that are physically allowed (Section 3.2.2) can be well-constrained. More details about the NG15yr dataset can be found in G. Agazie et al. (2023).
3 Methods
@@
A. Leleu et al. (2015, 2017) showed that a star-exoplanet-exotrojan system is largely degenerate with an ordinary star-exoplanet system. However, the following subsections discuss two ways to break this degeneracy in the context of pulsar binary systems. We ignore $\mathcal{O}(e_{\rm o})$ effects throughout this study, because the highest eccentricity of the pulsars studied is of the order $10^{-4}$.
3.1 Conjunction Inconsistency

### v2

# Searching for Exotrojans in Pulsar Systems

@@ Abstract
[…]
@@
facilities: GTC, CHIME, GBT, Arecibosoftware: PINT (Luo et al., 2019, 2021), rebound (Rein and Liu, 2012; Rein and Spiegel, 2015), ENTERPRISE (Ellis et al., 2019; Johnson et al., 2023), enterprise_extensions (Taylor et al., 2021; Johnson et al., 2023), PTMCMCSampler (Ellis and van Haasteren, 2017; Johnson et al., 2023), numpy (Harris et al., 2020), matplotlib (Hunter, 2007), corner (Foreman-Mackey, 2016).
I Introduction
Trojans are objects that reside in the $L_{4}$ or $L_{5}$ Lagrange points of two massive bodies. In our case, the relevant configuration consists of a pulsar, its low-mass companion (which we call the secondary), and a possible Trojan. The Trojan shares the secondary’s mean motion and trails or leads the secondary by $60^{\circ}$ throughout its orbit. In the Jupiter-Sun system alone, there are over ten thousand confirmed Trojans (e.g., Li et al., 2023) along with estimates of hundreds of thousands more that have diameters above $\sim$ 2 km (see, e.g., Jewitt et al., 2000; Yoshida and Nakamura, 2005). Every Solar System planet currently has Trojans of its own, barring Mercury—which may lack confirmed Trojans due to observational difficulties—and there are even Trojans of the Saturn-Tethys system.
The abundance of Trojans in our Solar System suggests that extrasolar Trojans (exotrojans) should also be common. Indeed, any robust Solar System formation theory must allow for their creation, and many of these theories do not just predict Trojan asteroids but also, surprisingly, Trojans of planetary mass (e.g., Laughlin and Chambers, 2002; Cresswell and Nelson, 2008; Lyra et al., 2009; Montesinos et al., 2020). Once formed, it has been shown that these Trojans can find themselves in relatively large stable regions (Danby, 1964; Giuppone et al., 2010; Leleu et al., 2015). In the case of circular orbits, the triangular Lagrange points are only stable if
@@ I Introduction
where $m_{\rm p}$ is the mass of the host star (a pulsar in this study), $m_{\rm c}$ is the mass of the secondary (the pulsar’s dominant companion), and $m_{\rm T}$ is the mass of the Trojan (the pulsar’s lowest mass companion; Gascheau 1843). Danby (1964) showed that introducing a non-zero orbital eccentricity generally decreases the window of stable masses.
@@
Unless the Trojan is librating around its stable point, the measured gravitational effects on the star from a single orbiting companion are degenerate with those of two co-orbiting companions up to $\mathcal{O}(e_{\rm{o}}^{2})$ (Leleu et al., 2017), where $e_{\rm o}$ is the orbital eccentricity. This degeneracy helps explain why, despite their prevalence in our Solar System and their accompanying formation models, no Trojans outside the Solar System have yet been confirmed. Ford and Gaudi (2006) showed that combining radial velocity data with transit-timing data breaks this degeneracy. Leleu et al. (2017) further formalized their approach into the robust $\alpha$-test method to include the results of several transits. Balsalobre-Ruza et al. (2024) applied this technique to identify one of the strongest candidates to date with a $3$-$\sigma$ detection, building on the earlier works of Lillo-Box et al. (2018a, b). These works are part of, or helped form the foundation for, the broader TROY project,111www.troy-project.com which seeks to discover the first extrasolar Trojan planets. While most of their work has been confined to optical observations of main-sequence stars, Balsalobre-Ruza et al. (2023) potentially identified an accumulation of gas and dust that might be a forming Trojan in the PDS 70 protoplanetary disk (e.g., Keppler et al., 2018; Haffert et al., 2019).
@@
Other studies have searched for photometric transits of the Trojans themselves. Janson (2013) used the Kepler space telescope to rule out transiting Trojans down to 1 R in 2244 systems, and Moldovan et al. (2010) used the MOST satellite to place a lunar mass upper limit on a Trojan in the HD 209458 system. Rather than looking for transits from individual Trojans, Hippke and Angerhausen (2015) looked for a cumulative Trojan effect across nearly 4000 planets by adding their Kepler light curves coherently by orbital phase, and identified potential transits of the stacked light curves at the expected $L_{4}$ and $L_{5}$ orbital phases. The main problem with identifying Trojans transiting their host stars is the fact that we expect exotrojans to have a broad range of orbital inclinations, as is the case for the Jupiter Trojans (Yoshida and Nakamura, 2005; Nesvornỳ et al., 2013; Bottke et al., 2023). Therefore, the dominant Kepler planet transiting its host star does not guarantee that any Trojan planets will also transit during any given observation. In contrast, the $\alpha$-test method does not rely on in-plane Trojans.
@@
While almost every exotrojan search has been done around main-sequence stars, as detailed above, no study has looked for them around pulsar binary systems. Just as we expect Trojans to form around ordinary stars during the protoplanetary disk phase (e.g., Montesinos et al., 2020), Trojans may also form around pulsars when a disk is present and then settle via gas drag (Lyra et al., 2009) into stable so-called tadpole (T-type) orbits. Such a disk can form in at least two possible ways. One type of pulsar disk, the supernova fallback disk, forms when some of the material from the supernova that formed the pulsar returns to the pulsar and collapses into a disk (Behrens et al., 2020; Xu et al., 2024). Such disks are a plausible formation channel for pulsar planets (Lin et al., 1991; Wolszczan and Frail, 1992), and would explain why the planets around PSR B1257+12 have similar inclinations (Konacki and Wolszczan, 2003).
The other possible type of pulsar disk forms when the compact pulsar in a binary system strips material from its evolving, ablated companion. This material forms an accretion disk, which collects onto the pulsar and transitions the pulsar into a millisecond pulsar (MSP) (Archibald et al., 2009; Papitto and Martino, 2021). Before the accretion disk dissipates, it may form planets (Patruno and Kama, 2017; Behrens et al., 2020), an asteroid belt (Shannon et al., 2013), or Trojan planets in 1:1 resonance with the pulsar’s companion, as in the case of protoplanetary-disk-formed Trojans (Lyra et al., 2009; Montesinos et al., 2020). Potential Trojans, too, may be stripped of their material, but the timescale of complete dissipation is unclear, with some estimates ranging up to a few Gyr for the secondary (e.g. Ginzburg and Quataert, 2020). This picture is complicated by the fact that material from the secondary could continue to fall into the Triangular Lagrange points. In any case, detecting an interloping Trojan may be possible even if it lasts just a few hundred Myr before complete dissipation.
The above pulsar Trojan formation discussion lays out plausible channels, but the possible existence of pulsar Trojans remains speculative. Future theoretical work is required to draw definitive connections between proto-planetary disks and pulsar-centered disks with respect to Trojan creation. Although the existence of pulsar Trojans is uncertain, the history of serendipitous astronomical discoveries in the last several decades—including pulsars, fast radio bursts, and hot Jupiters—motivates their search. In any case, our results provide an independent observational baseline for future theoretical studies, such as hydrodynamical simulations, of pulsar Trojan viability.
@@
The first exoplanet was discovered around a pulsar (Wolszczan and Frail, 1992), and ever since there has been a growing interest in discovering massive bodies in the environs of pulsars. Besides studies looking for planets in the ordinary P-type orbit (e.g., Bailes et al., 2011; Spiewak et al., 2018; Behrens et al., 2020; Niţu et al., 2022), pulsar asteroid belts (Shannon et al., 2013; Jennings et al., 2020b) and freely-floating planets gravitationally scattering with pulsars (Jennings et al., 2020a; Dey et al., 2026) have also been investigated. In this work, we turn the search to Trojans that may reside in black widow pulsar systems—binaries characterized by an MSP with a low-mass ($\sim 0.01\,\mathrm{M_{\odot}}$) companion (e.g., Shahbaz et al., 2017)—and are therefore likely to follow the stability criterion given in Equation 1.
@@
To conduct the first-ever search for Trojan planets in pulsar binary systems, we use the North American Nanohertz Observatory for Gravitational Waves (NANOGrav) 15-year (NG15yr) narrow-band dataset (Agazie et al., 2023) as well as optical-band light curves of PSR J1641+8049 from Kirichenko et al. (2024) to search for pulsar systems bearing Trojans in the T-type configuration. We ignore horseshoe configurations because these systems do not yield a closed-form waveform in the pulse times of arrival (TOAs; see Section II.2 and Leleu et al. 2015). A Trojan in the T-type configuration librates around a particular triangular Lagrange point, while a Trojan in the horseshoe configuration will repeatedly librate between $L_{4}$ and $L_{5}$ going through $L_{3}$. Excluding horseshoe orbiters does not greatly limit our search, as not only are they less common in the Solar System (e.g., Kaplan and Cengiz, 2020), but also their stability criterion are stricter (e.g., Laughlin and Chambers, 2002).
We outline our methods in Section II, select our observed data in Section III, display and discuss our results in Section IV, and conclude with a broader discussion in Section V.
II Methods
@@
Leleu et al. (2015, 2017) showed that the signal observed from a star-exoplanet-exotrojan system is largely degenerate with that from an ordinary star-exoplanet system. This is due to the secondary and Trojan forming a combined effective companion. The TOAs will indicate the pulsar and a combined effective companion are orbiting a center of mass. The following subsections discuss two ways to break this degeneracy in the context of pulsar binary systems.
The first method uses both optical light curves and radio pulsar TOAs. Both the light curves and TOAs can independently measure the time of superior conjunction (TSC), which is when the pulsar’s companion is furthest from the Earth in its orbit. A discrepancy between the light-curve-measured TSC and TOA-measured TSC can reveal the presence of a mass accumulation along the orbit in the system. This is an adaptation the method developed by Ford and Gaudi (2006) for exoplanets orbiting main-sequence stars. In their case, they compare the transiting time of the secondary measured from photometry—in other words, the time of inferior conjunction (TIC)—with the estimated transiting time from radial velocities. The orbital planes of the secondaries in our study are not “edge-on” with respect to our LOS, so we cannot measure their direct transits. However, we can measure their TSCs which corresponds to peak optical flux, as the secondaries are tidally locked with pulsar. Note that we assume negligible flux contributions from the pulsar and any Trojans. Analogously to radial velocities, TOAs trace the Keplerian orbit from the effective orbiting mass (secondary-exotrojan) by measuring the pulsar LOS position. Therefore, the TOAs uniquely determine the relevant Keplerian parameters (e.g., Taylor et al., 2024), which give the TSC. See Section II.1 for more details and Figures 1A and 1C of Ford and Gaudi 2006 for an illustration of the concept.
The second method requires only TOAs, but only works if any Trojan and secondary in the system is librating around the stable Lagrange point. If the Trojan and secondary are librating with some frequency, the effective companion-pulsar center of mass will also oscillate at the same frequency. A pulsar orbiting an oscillating center of mass no longer produces TOAs degenerate with those of a two-body system, and allows us to measure important properties and masses of the system. This method comes directly from Leleu et al. (2015), where they derive both the line-of-sight position and radial velocity of a main-sequence star that hosts librating co-orbitals. As mentioned above, we can use the TOAs to measure the pulsar’s line-of-sight (LOS) position. We expand upon their LOS position derivation in Appendix A, including a minor algebraic correction. This libration-dependent method can be used to detect a single Trojan in a pulsar binary system. In the case of a Trojan swarm, this method remains effective in detecting the highest-mass Trojan, provided this Trojan is massive enough to allow gravitational effects from the other Trojans to be ignored. See Section II.2 for more details.
In this work, we quote posterior medians with uncertainties enclosing the 68% credible interval. We quote our upper limits at 95% credibility. We use both the parentheses-on-the-last-digit(s) convention and the $\pm$ notation to express these 68% uncertainties. Throughout, $\log$ denotes the base-10 logarithm.
II.1 Conjunction Inconsistency